Please help with this question.



Solve the equation for x:
3(x-4) = 4 - 2(x+2)

Answer by stanbon(75887)

Solve the equation for x:
3(x-4) = 4 - 2(x+2)
----
3x-12 = 4 -2x -4
5x = 12
x = 12/5
-------------
Cheers,
Stan H.
=============